OpenStudy (anonymous):

9 = -3 + n + 2n (&please show your steps)

OpenStudy (anonymous):

9 = -3 + n + 2n 12=n+2n 12=3n 12/3=3n/3 4=n

OpenStudy (anonymous):

9 + 3 = 3n 12 = 3n 12/3 = n

OpenStudy (anonymous):

i need help with cross reducing anyone know?